a. If Canace Company, with a break-even point at $960,000 of sales, has actual sales of $1,200,000, what is the margin of safety expressed (1) in dollars and (2) as a percentage of sales? Round the percentage to the nearest whole number.

1. $240,000

2. 20%

b. If the margin of safety for Canace Company was 20%, fixed costs were $1,875,000, and variable costs were 80% of sales, what was the amount of actual sales (dollars)?
(Hint: Determine the break-even in sales dollars first.)
$
Answer #1
Concepts and reason

CVP analysis: CVP is one of the techniques of decision accounting to achieve the targeted profits by changing different variables. The relationship between cost, profit and sales volume provides the basis for the manager to take effective steps about the future profits.

Fixed Cost: Fixed cost is a cost that remains same irrespective of the increase or decrease in the value of goods or any services rendered. It is the cost paid by the company that does not depend on the activities concerned with the business.

Variable Cost: Variable cost is a cost that varies according to the output produced or any service rendered. It is the cost paid by the company that depends on the activities concerned with the business.

Fundamentals

Selling price: Selling price is the price at which products is sold in the market. It includes the total cost incurred to the product and profit.

Breakeven point: The point of sales at which there is no profit or loss is called as Break Even Point (BEP). It means the total cost equals the total revenue at this point.

Margin of safety: Margin of safety is an excess of budgeted sales than break-even sales.

Margin of safety ratio: Margin of safety ratio is calculated by dividing margin of safety by expected sales. It is also called as margin of safety as a percentage of sales.

Breakeven point sales in dollars: The point, which represents sales dollars where total revenue equals total costs, is referred to as breakeven point in sales. The following formula is used for calculating breakeven point sales in dollars:

Breakevensalesindollars=FixedcostsContributionmarginratio{\rm{Break - even sales in dollars = }}\frac{{{\rm{Fixed costs}}}}{{{\rm{Contribution margin ratio}}}}

a.

1)

Calculate margin of safety in dollars.

Marginofsafetyindollars=(TotalactualsalesindollarsBreakevenpointindollars)=$1,200,000$960,000=$240,000\begin{array}{c}\\{\rm{Margin of safety in dollars = }}\left( \begin{array}{l}\\{\rm{Total actual sales in dollars }}\\\\ - {\rm{ Break - even point in dollars}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = \$ 1,200,000 }} - {\rm{ \$ 960,000 }}\\\\{\rm{ = \$ 240,000}}\\\end{array}

Therefore, the margin of safety in dollars is $240,000.

a)

2)

Calculate margin of safety percentage.

Marginofsafetypercentage=(MarginofsafetyTotalactualsales)×100=($240,000$1,200,000)×100=20%\begin{array}{c}\\{\rm{Margin of safety percentage = }}\left( {\frac{{{\rm{Margin of safety}}}}{{{\rm{Total actual sales}}}}} \right)\; \times \;{\rm{100 }}\\\\{\rm{ = }}\left( {\frac{{{\rm{\$ 240,000}}}}{{{\rm{\$ 1,200,000}}}}} \right)\; \times \;{\rm{100 }}\\\\{\rm{ = 20\% }}\\\end{array}

Therefore, the margin of safety percentage is 20%.

b)

Calculate amount of actual sales in dollars.

Amountofactualsalesindollars}=BreakevensalesindollarsBreakevensalespercentage=$9,375,00080%=$11,718,750\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Amount of actual}}\\\\{\rm{ sales in dollars}}\\\end{array} \right\}{\rm{ = }}\frac{{{\rm{Break - even sales in dollars}}}}{{{\rm{Break - even sales percentage}}}}{\rm{ }}\\\\{\rm{ = }}\frac{{{\rm{\$ 9,375,000}}}}{{{\rm{80\% }}}}{\rm{ }}\\\\{\rm{ = \$ 11,718,750}}\\\end{array}

Therefore, the amount of actual sales in dollars is $11,718,750.

Working note:

Calculate break-even sales percentage.

Breakevensalespercentage=(TotalsalespercentageMarginofsafetypercentage)=100%20%=80%\begin{array}{c}\\{\rm{Break - even sales percentage = }}\left( \begin{array}{l}\\{\rm{Total sales percentage }}\\\\ - {\rm{ Margin of safety percentage}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = 100\% }} - {\rm{ 20\% }}\\\\{\rm{ = 80\% }}\\\end{array}

Calculate contribution margin percentage.

Contributionmarginpercentage=(TotalsalespercentageVariablecostpercentage)=100%80%=20%\begin{array}{c}\\{\rm{Contribution margin percentage = }}\left( \begin{array}{l}\\{\rm{Total sales percentage }}\\\\ - {\rm{ Variable cost percentage}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = 100\% }} - {\rm{ 80\% }}\\\\{\rm{ = 20\% }}\\\end{array}

Calculate break-even sales in dollars.

Breakevensalesindollars=FixedcostsContributionmarginratio=$1,875,00020%=$9,375,000\begin{array}{c}\\{\rm{Break - even sales in dollars = }}\frac{{{\rm{Fixed costs}}}}{{{\rm{Contribution margin ratio}}}}{\rm{ }}\\\\{\rm{ = }}\frac{{{\rm{\$ 1,875,000}}}}{{{\rm{20\% }}}}{\rm{ }}\\\\{\rm{ = \$ 9,375,000}}\\\end{array}

Ans: Part a.1

Margin of safety in dollars is $240,000.

Part a.2

Margin of safety percentage is 20%.

Part b

Amount of actual sales in dollars is $11,718,750.